2.You decide to start saving for your retirement, in 25 years time. Today you make an initial lump sum payment of 10,000, then decide to save 500 per semester and expect an average return of 6.6%(comp.semesterly or semiannually). How much will you have in the end, assuming you pay the money in the beginning of each semester? 3.Your bank has just launched a savings scheme which pays an interest at 5.15% monthly compounded, over 10 years. If you invest 100 at the end of each month, how much will you have at the end of those 10 years?
Answer #1

2.

Amount at the end of 25 years

[10000 * (1 + (0.066/2))^(25*2)] + [500 * Future value annuity due factor @ 3.3% for 50 periods]

= [10000 * (1+0.033)^50] + [500 * 127.41]

= [10000 * 5.07] + 63705

= 50700 + 63705

= 114,405.

3.

Interest Rate per annum = 5.15%

Interest Rate per month = 5.15/12

= 0.43%

Investment made per month = 100

Amount after 10 years = 100 * (Future value annuity factor @ 0.43% for 120 periods)

= 100 * 156.62

= 15,662.